OpenStudy (anonymous):

Sheila went to the store and bought 28 apples. She bought twice as many red apples as she bought green, and she bought one-fourth as many yellow as she did red. How many of each type of apple did she buy? A. 16 red; 8 green; 4 yellow B. 12 red; 13 green; 3 yellow C. 14 red; 7 green; 7 yellow D. 18 red; 8 green; 2 yellow

OpenStudy (anonymous):

the answer is A :)

OpenStudy (anonymous):

Thanks. I got it correct on my lesson. Thanks so much

OpenStudy (anonymous):

2x + x + (1/4)2x = 28 3x + x/2 = 28 (6x + x)/2 = 28 7x = 56 x = 8 = green apples

OpenStudy (anonymous):

Just so you know how to do the problem.

OpenStudy (anonymous):

thanks.